# Unsourced Adversarial CAPTCHA: A Bi-Phase Adversarial CAPTCHA Framework

### v2

@@ Abstract
Traditional CAPTCHA schemes are increasingly vulnerable to automated attacks powered by deep neural networks (DNNs). Existing adversarial attack methods often rely on original image characteristics, resulting in distortions that hinder human interpretation and limiting their applicability in scenarios with no initial input images. To address these challenges, we propose the Unsourced Adversarial CAPTCHA (UAC), a novel framework generating high-fidelity adversarial examples guided by attacker-specified text prompts. Leveraging a Large Language Model (LLM), UAC enhances CAPTCHA diversity and supports both targeted and untargeted attacks. For targeted attacks, the EDICT method optimizes dual latent variables in a diffusion model for superior image quality. In untargeted attacks, especially for black-box scenarios, we introduce bi-path unsourced adversarial CAPTCHA (BP-UAC), a two-step optimization strategy employing multimodal gradients and bi-path optimization for efficient misclassification. Experiments show BP-UAC achieves high attack success rates across diverse systems, generating natural CAPTCHAs indistinguishable to humans and DNNs.
@@ I Introduction
CAPTCHA (Completely Automated Public Turing test to tell Computers and Humans Apart) is a foundational cybersecurity mechanism. Its core function is to distinguish legitimate human users from automated bots. This technology presents specific computational challenges that are easily solvable by humans but difficult for machines. Common implementations can be categorized into text-based CAPTCHAs e.g., interpreting distorted alphanumeric sequences and solving arithmetic problems, image-based CAPTCHAs e.g., recognizing objects in image grids and completing slider puzzles, and audio-based CAPTCHAs. CAPTCHA systems establish critical barriers against automated abuse. They are widely deployed to prevent malicious activities, such as credential stuffing, spam registrations, ticket scalping, data scraping, and comment spam. By filtering out automated attacks, these mechanisms safeguard digital services while maintaining system integrity.
@@
As AI systems become more proficient at tasks traditionally requiring human intelligence, they also pose unintended challenges in areas where maintaining human control and security is critical. For instance, image recognition technologies based on DNN model have grown highly effective at deciphering even the most complex CAPTCHA systems, which are widely used to prevent automated bots and malicious activities. CAPTCHAs, originally designed to differentiate between humans and machines, are now increasingly vulnerable to attacks powered by advanced deep learning models. These models can break CAPTCHA defenses with remarkable precision, rendering conventional CAPTCHA systems less effective at safeguarding online platforms against automated attacks.
@@
Figure 2: Our proposed bi-path unsourced adversarial CAPTCHA (BP-UAC) attack framework
@@
Adversarial attacks, which exploit the sensitivity of deep learning models to specific, carefully designed perturbations, have emerged as a powerful tool to evaluate model robustness [1, 2]. These attacks introduce subtle, targeted noise to mislead models into incorrect predictions, serving as an effective tool for evaluating the robustness of AI systems. This concept has been creatively extended to adversarial CAPTCHAs, where adversarial perturbations are used to enhance CAPTCHA robustness against automated recognition. This novel approach has started gaining traction in recent years, particularly among leading AI enterprises, and represents an innovative shift in security strategies for CAPTCHA systems [3, 4].
Although adversarial CAPTCHAs represent a significant advancement in securing systems against automated attacks, adversarial CAPTCHA systems still face several challenges. First, the limited variety of CAPTCHA images, often constrained by issues like copyright, restricts the diversity of generated adversarial examples, which in turn can inflate the cost of model training. Second, to effectively resist recognition attempts by various black-box models, adversarial CAPTCHAs require strong generalization abilities, often achieved by amplifying the intensity of adversarial perturbations. However, increasing the strength of these perturbations can introduce more visual noise, creating difficulties for legitimate users during the verification process and adversely impacting the usability and overall user experience of CAPTCHA systems. The potential of diffusion models offers a promising direction for further enhancing CAPTCHA security and usability, ensuring robustness against an even wider array of attack methodologies.
@@
To address the above issues, we propose the unsourced adversarial CAPTCHA (UAC) attack framework in the white-box scenario. Specifically, before initiating the attack, we use a large language model to convert the attacker’s input prompt into a concise and clear sentence, thereby ensuring the accuracy of the generated image. Unlike other adversarial attack methods based on diffusion model, we adopt EDICT’s [5] dual latent variable optimization approach. This method approximates the initial input by performing one-step denoising at each time step, allowing the model gradients to propagate backward through the entire chain and resolving the instability and error accumulation issues inherent in adversarial attack methods based on diffusion model. Through the precise inversion process, UAC effectively avoids error propagation seen in traditional methods, ensuring the stability and high quality of adversarial example generation. In the more challenging and practical black-box attack scenario, we further introduce the bi-path unsourced adversarial CAPTCHA (BP-UAC). This method enhances optimization efficiency and attack performance by integrating multi-model gradient optimization and proposing a novel bi-path optimization strategy during the attack process. It ensures that the generated adversarial examples exhibit high transferability in black-box environments.
@@
$\bullet$ We propose the first unsourced attack framework based on text-guided generation of adversarial examples, which enables the generation of adversarial examples during the diffusion process instead of relying on the original input images. The method effectively improves the diversity of generated adversarial examples and solves the optimization problem of adversarial example perturbation concealment, significantly improving the search space for adversarial attacks.
$\bullet$ We propose different attack frameworks for the white-box and black-box scenarios. In the white-box scenario, we propose the unsourced adversarial CAPTCHA (UAC) framework, which uses a large language model to convert the attacker’s input into a clear sentence, ensuring accurate image generation. We also employ EDICT’s dual latent variable optimization to address instability and error propagation in other adversarial attack methods based on diffusion model, ensuring stable and high-quality adversarial examples. In the black-box scenario, we further propose BP-UAC, which enhances optimization efficiency by integrating multi-model gradient optimization and employing a bi-path optimization strategy, improving the transferability and robustness of adversarial examples.
$\bullet$ Extensive experiments have demonstrated that our approach can guarantee high attack success rate (ASR) for deep neural network models with different structures under different white- and black-box scenarios while generating unsourced Adversarial CAPTCHAs that are indistinguishable from clean examples.
To the best of our knowledge, we are the first to generate unsourced adversarial CAPTCHA using the diffusion model. At the same time, by means of a novel bi-path optimization strategy, we achieve for the first time a near 100% ASR against an unknown black-box model when generating adversarial examples using the generative model. This opens up new possibilities for adversarial research.
@@ II RELATED WORK
In recent years, there have been various innovative approaches to adversarial image attacks. MUTEN [6] enhances the success rate and robustness of gradient-based adversarial attacks by utilizing diverse variant models. GADT [7] improves the migrability of adversarial examples by optimizing the data enhancement parameters, which is particularly suitable for black-box and query attacks. MGAA [8] improves the mobility of attacks using meta-learning methods and enhances the success rate of attacks by narrowing the difference in gradient directions in white-box and black-box environments. U-GAN [9] constructs unconstrained adversarial exexample through GAN networks, breaking through the restriction on the perturbation range of traditional gradient-based attack methods, making the generated adversarial example more aggressive and effective in bypassing many existing defense mechanisms. AdvDiff [10] generates unconstrained adversarial example by utilizing the denoising process of the diffusion model, demonstrating the potential of this model in adversarial attacks. DiffAttack [11] proposes an attack strategy specifically for countering the purification defense of the diffusion model, which successfully generates adversarial example by bypassing the purification process, revealing potential loopholes in the defense mechanisms of the diffusion model and driving new challenges in the field of adversarial attacks.
@@
Not only in the field of image recognition, but also in the field of security, researchers have similarly tried to incorporate undetectable perturbations into CAPTCHAs to fight against machine intrusion. Shi et al. [12] proposed an aCAPTCHA system, which enhances the security of ordinary CAPTCHAs by generating adversarial examples. This approach makes it difficult for deep learning models to recognize them by adding adversarial perturbations to the images, while still allowing human users to pass normally. Wen et al. [13] explored methods to generate stronger CAPTCHA challenges through adversarial attacks such as Iterative FGSM (I-FGSM) [14] and DeepFool [15]. Their method, which focuses on perturbation processing, improves the resistance of these visual challenges to machine learning models, making them harder for automated systems to crack. Zhang et al. [17] explored the application of adversarial examples on different image classes of CAPTCHAs, and investigated how to improve the robustness of image CAPTCHAs by using adversarial methods by analyzing the effect of adversarial example perturbation on it.
While all of the above methods optimize the perturbations at different levels, it is always necessary to generate image adversarial examples based on the corresponding benign images, and the fundamental flaws of the adversarial attack methods in this regard limit the search space of the perturbations and the high intensity of the attack performance can make the perturbations too noisy, which affects their visualization by humans. In contrast to the above work, our focus is on introducing gradient information of DNNs in the diffusion generation process and enabling their stable integration into the diffusion model to generate high-quality adversarial samples.
@@ III-A Traditional adversarial scenarios
Traditional adversarial attacks can be categorized into untargeted and targeted attacks. To cope with different application scenarios, our approach discusses targeted and untargeted attacks, respectively. Targeted attacks aim to shift the model’s predictions to a targeted category specified by the attacker. This attack can be considered “targeted” because the attacker wants the model to output a specific mislabel when confronted with a modified example. Specifically, a traditional targeting attack is formulated as follows :
$\begin{gathered}x^{\prime}=\arg\min_{x^{\prime}}\ell(f(x^{\prime}),y_{\mathrm{target}})\\ \begin{aligned} \mathrm{s.t.}\>x^{\prime}=x+\delta\end{aligned}\end{gathered}$ (1)
Instead, the untargeted attack aims to make the prediction results different from the original category without specifying the wrong category. The attacker only needs the model to output any of the wrong categories when it sees the antagonistic example.
$\begin{gathered}x^{\prime}=\arg\max_{x^{\prime}}\ell(f(x^{\prime}),y_{\mathrm{target}})\\ \begin{aligned} \mathrm{s.t.}\>x^{\prime}=x+\delta\end{aligned}\end{gathered}$ (2)
From Eq. 1 and Eq. 2, it can be concluded that traditional adversarial attacks mainly rely on specify benign images $x$ as the basis for generating adversarial examples $x^{\prime}$.
@@ III-B Diffusion model
Denoising Diffusion Models (DDMs) [18] are a type of generative model that produces images by progressively removing noise, starting from pure noise and iterating until a clean image is generated. The process consists of two main stages: forward diffusion and reverse diffusion. Forward Diffusion Process: In this stage, clean images are corrupted by adding Gaussian noise in a series of steps, ultimately transforming the image into pure noise. At each step $t$, the noisy image $x_{t}$ is computed as:
$x_{t+1}=\sqrt{a_{t}x_{t}+\sqrt{1-\alpha_{t}\epsilon}}$ (3)
where is $a_{t}$ noise-scaling factor and $\epsilon$ represents Gaussian noise.
Reverse Diffusion Process: Starting from pure noise, the model denoises the image iteratively, generating a series of intermediate images that gradually resemble the original image. The reverse process is modeled by:
$p_{\theta}\left(x_{t-1}\mid x_{t}\right)$ $=\mathcal{N}\left(x_{t-1};\mu_{\theta}\left(x_{t},t\right),\Sigma_{\theta}\left(x_{t},t\right)\right)$ (4)
where $\mu_{\theta}$ and $\Sigma_{\theta}$ are the learned parameters for mean and variance, guiding the denoising process.
To accelerate sampling, Denoising Diffusion Implicit Models (DDIM) [19] introduce a non-stochastic, efficient sampling approach. DDIM avoids fully stochastic sampling, offering a faster alternative by approximating the reverse steps with an implicit formula:
$x_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\cdot\mu_{\theta}(x_{t},t)+\sqrt{1-\bar{\alpha}_{t-1}}\cdot\epsilon$ (5)
This formula reduces the number of steps required to generate an example, making the model more practical for real-time applications.
@@ III-C EDICT
In recent years, EDICT [5] has been proposed to address the problem of reconstruction and content loss in model-generated images due to error propagation during DDM [20] backpropagation. EDICT mathematically and precisely inverts real and model-generated images by coupling latent noise variables. Specifically, EDICT utilizes two alternating sequences, $x_{t}$ and $y_{t}$, in the symmetric coupling layer of the DDM. These sequences are used to track and reconstruct the states of the latent variables during the generation process. The forward updating of $x_{t}$ and $y_{t}$ during DDM inverse propagation is then realized through the following coupling steps:
$x_{t}^{inter}$ $=a_{t}\cdot x_{t}+b_{t}\cdot\Theta_{(t,C)}(y_{t})$ (6)
$y_{t}^{inter}$ $=a_{t}\cdot y_{t}+b_{t}\cdot\Theta_{(t,C)}(x_{t}^{inter})$
$x_{t-1}$ $=p\cdot x_{t}^{inter}+(1-p)\cdot y_{t}^{inter}$
$y_{t-1}$ $=p\cdot y_{t}^{inter}+(1-p)\cdot x_{t-1}$
where $a_{t}$ and $b_{t}$ denote the coefficients from time to time, EDICT includes the internal transformations and backward updates of the variables. It generates a new $x_{t-1}$ and $y_{t-1}$ by mixing the internal representations of $x_{t}$ and $y_{t}$ through the averaging parameter $p\in(0,1)$, which is derived using linear equations through the above equation to ensure that the entire diffusion process can be performed exactly in reverse:
@@
$x_{t+1}$ $=(x_{t+1}^{inter}-b_{t+1}\cdot\Theta_{(t+1,C)}(y_{t+1}))/a_{t+1}$

### v3

# Defensive Adversarial CAPTCHA: A Semantics-Driven Framework for Natural Adversarial Example Generation

@@ Abstract
Traditional CAPTCHA (Completely Automated Public Turing Test to Tell Computers and Humans Apart) schemes are increasingly vulnerable to automated attacks powered by deep neural networks (DNNs). Existing adversarial attack methods often rely on the original image characteristics, resulting in distortions that hinder human interpretation and limit their applicability in scenarios where no initial input images are available. To address these challenges, we propose the Unsourced Adversarial CAPTCHA (DAC), a novel framework that generates high-fidelity adversarial examples guided by attacker-specified semantics information. Leveraging a Large Language Model (LLM), DAC enhances CAPTCHA diversity and enriches the semantic information. To address various application scenarios, we examine the white-box targeted attack scenario and the black-box untargeted attack scenario. For target attacks, we introduce two latent noise variables that are alternately guided in the diffusion step to achieve robust inversion. The synergy between gradient guidance and latent variable optimization achieved in this way ensures that the generated adversarial examples not only accurately align with the target conditions but also achieve optimal performance in terms of distributional consistency and attack effectiveness. In untargeted attacks, especially for black-box scenarios, we introduce bi-path unsourced adversarial CAPTCHA (BP-DAC), a two-step optimization strategy employing multimodal gradients and bi-path optimization for efficient misclassification. Experiments show that the defensive adversarial CAPTCHA generated by BP-DAC is able to defend against most of the unknown models, and the generated CAPTCHA is indistinguishable to both humans and DNNs.
@@ I Introduction
CAPTCHA (Completely Automated Public Turing Test to Tell Computers and Humans Apart) is a fundamental cybersecurity mechanism. Its core function is to distinguish legitimate human users from automated bots. This technology presents specific computational challenges that are easily solvable by humans but difficult for machines. By filtering out automated attacks, these mechanisms safeguard digital services while maintaining system integrity.
@@
Generally speaking, the effectiveness of a CAPTCHA can be evaluated based on its security. This refers to the CAPTCHA’s ability to defend against unknown hacking techniques used by attackers, as well as its user-friendliness. However, with the rapid development of artificial intelligence [1, 2, 3, 4, 5, 6, 7], DNN models have achieved remarkable results in fields such as image recognition and image segmentation [8, 9, 10, 11, 12, 13]. They also pose unintended challenges in areas where maintaining human control and security is critical. For instance, image recognition technologies based on DNN models have grown highly effective at deciphering even the most complex CAPTCHA systems. CAPTCHAs, originally designed to differentiate between humans and machines, are now increasingly vulnerable to attacks powered by advanced deep learning models. These models can break CAPTCHA defenses with remarkable precision, rendering conventional CAPTCHA systems less effective at safeguarding online platforms against automated attacks.
@@
Figure 2: Our proposed bi-path unsourced adversarial CAPTCHA (BP-DAC) attack framework.
@@
Adversarial attacks [14, 15, 16], which exploit the sensitivity of deep learning models to specific and carefully designed perturbations, have emerged as a powerful tool to evaluate model robustness [17, 18]. These attacks introduce subtle, targeted noise to mislead models into incorrect predictions, serving as an effective tool for evaluating the robustness of AI systems. This concept has been creatively extended to adversarial CAPTCHAs, where adversarial perturbations are used to enhance CAPTCHA robustness against automated recognition. This novel approach has started to gain traction in recent years, particularly among leading AI companies, and represents an innovative change in security strategies for CAPTCHA systems [19, 20].
Although adversarial CAPTCHAs represent a significant advancement in system security against automated attacks, adversarial CAPTCHA systems still face numerous challenges. The limited availability of image types (due to copyright restrictions, etc.) results in insufficient diversity of adversarial samples. Additionally, while increasing perturbation intensity to ensure strong defense against black-box models, this approach compromises visual naturalness, thereby harming user experience and usability. Furthermore, generating CAPTCHAs that are both natural and widely resistant to DNN attacks requires a vast search space, which traditional pixel-level perturbation methods struggle to handle. Diffusion models offer a new direction for enhancing both security and usability, as their potential to unlock search space enables the generation of more natural yet adversarial images, thereby strengthening robustness. However, existing diffusion model-based attack techniques still rely on original samples, limiting the full realization of their adversarial strength.
@@
To address these issues, we designed the DAC framework. By shifting input information from the pixel-level to the semantic level, we overcame the limitations of traditional adversarial sample generation methods, which are constrained by the pixel search space. This enables defensive CAPTCHA designers to generate adversarial CAPTCHA images rich in semantic information simply by inputting the desired image categories. This reduces the need for data collection and manual annotation of verification images. Specifically, before the attack, in the face of white-box scenarios that are easier to defend against, we utilize LLM to convert the categories entered by CAPTCHA designers into more complex semantic information, thereby ensuring the accuracy of the generated images. Additionally, unlike other adversarial attack methods based on diffusion models, we design parameter-shared coupled noise vector pairs during the diffusion model generation process, achieving robust inversion through an iterative bidirectional calibration mechanism at discrete time steps. This method approximates the initial input by performing a denoising operation at each time step, allowing the model gradient to propagate backward through the entire chain. This solves the inherent instability and error accumulation problems of diffusion-based adversarial attack methods. It ensures the stability and high quality of adversarial example generation. In more challenging and practical black-box attack scenarios, we further propose the BP-DAC framework, which proposes a novel bi-path optimization strategy. This method enhances optimization efficiency and attack performance by further aggregating multi-model gradients during the attack optimization process while simultaneously guiding the generative model in the direction of the suboptimal and lowest category gradients of the target model.
@@
$\bullet$ To address the current issues with CAPTCHA systems, we propose the DAC framework, which innovatively transforms the input of CAPTCHA creators from clean existing images to semantic information. By incorporating EDICT into the process of generating adversarial samples using diffusion models, our method achieves high-fidelity adversarial sample generation through precise coupling inversion. Compared to existing methods, it generates more accurate images while reducing error accumulation, maintaining stealthiness, and without requiring additional training or data adaptation.
$\bullet$ When faced with unknown black-box models, which are more common in real-world applications, we further proposed the BP-DAC framework, which integrates the gradients of multiple models into the guidance process and simultaneously guides the attack toward the suboptimal category and lowest category gradient of the target model, thereby improving optimization efficiency and the defense capability of CAPTCHAs against unknown models.
$\bullet$ Extensive experiments have demonstrated that our approach can guarantee a high attack success rate (ASR) for deep neural network models with different structures under different white- and black-box scenarios while generating unsourced Adversarial CAPTCHAs that are indistinguishable from clean examples.
To the best of our knowledge, we are the first to generate an unsourced adversarial CAPTCHA using the diffusion model. At the same time, by means of a novel bi-path optimization strategy, we achieve for the first time a near 99.4% ASR against an unknown black-box model when generating adversarial examples using the generative model. This opens up new possibilities for adversarial research.
@@ II RELATED WORK
In recent years, various innovative approaches to adversarial image attacks have emerged. MUTEN [21] enhances the success rate and robustness of gradient-based adversarial attacks by utilizing diverse variant models. GADT [22] improves the migrability of adversarial examples by optimizing the data enhancement parameters, which is particularly suitable for black-box and query attacks. MGAA [23] improves the mobility of attacks using meta-learning methods and enhances the success rate of attacks by narrowing the difference in gradient directions in white-box and black-box environments. U-GAN [24] constructs unconstrained adversarial examples through GAN networks, breaking through the restriction on the perturbation range of traditional gradient-based attack methods, making the generated adversarial example more aggressive and effective in bypassing many existing defense mechanisms. AdvDiff [25] generates unconstrained adversarial examples by utilizing the denoising process of the diffusion model, demonstrating the potential of this model in adversarial attacks. DiffAttack [26] proposes an attack strategy specifically for countering the purification defense of the diffusion model, which successfully generates adversarial examples by bypassing the purification process, revealing potential loopholes in the defense mechanisms of the diffusion model and driving new challenges in the field of adversarial attacks.
@@
Not only in the field of image recognition, but also in the field of security, researchers have attempted to incorporate undetectable perturbations into CAPTCHAs to counter machine intrusion. Shi et al. [27] proposed an aCAPTCHA system, which enhances the security of ordinary CAPTCHAs by generating adversarial examples. This approach makes it difficult for deep learning models to recognize them by adding adversarial perturbations to the images, while still allowing human users to pass normally. Wen et al. [28] explored methods to generate stronger CAPTCHA challenges through adversarial attacks, such as Iterative FGSM (I-FGSM) [29] and DeepFool [30]. Their method, which focuses on perturbation processing, enhances the resistance of these visual challenges to machine learning models, making them more difficult for automated systems to crack. Zhang et al. [31] explored the application of adversarial examples on different image classes of CAPTCHAs, and investigated how to improve the robustness of image CAPTCHAs by using adversarial methods and analyzing the effect of adversarial example perturbation on it.
Although the above methods can generate natural adversarial examples using various techniques, they still require corresponding benign images to generate adversarial examples, or they cannot generate adversarial examples with complex semantic information. This is because the fundamental flaw of adversarial attack methods in this regard limits the search space of perturbations, and the high intensity of attack performance may lead to excessive perturbations, thereby affecting human visual perception. Unlike the above work, our approach breaks out of the traditional pixel space and generates complex adversarial examples based on intricate semantic information.
@@ III-A Traditional adversarial scenarios
Traditional adversarial attacks can be categorized into untargeted and targeted attacks. To cope with different application scenarios, our approach discusses both targeted and untargeted attacks. Targeted attacks aim to shift the model’s predictions to a targeted category specified by the attacker. This attack can be considered “targeted” because the attacker wants the model to output a specific mislabel when confronted with a modified example. Specifically, a traditional targeted attack is formulated as follows:
$\begin{gathered}x^{\prime}=\arg\min_{x^{\prime}}\ell(f(x^{\prime}),y_{\mathrm{target}})\\ \begin{aligned} \mathrm{s.t.}\>x^{\prime}=x+\delta.\end{aligned}\end{gathered}$ (1)
Instead, the untargeted attack aims to make the prediction results different from the original category without specifying the wrong category. The attacker only needs the model to output any of the wrong categories when it sees the antagonistic example:
$\begin{gathered}x^{\prime}=\arg\max_{x^{\prime}}\ell(f(x^{\prime}),y_{\mathrm{target}})\\ \begin{aligned} \mathrm{s.t.}\>x^{\prime}=x+\delta.\end{aligned}\end{gathered}$ (2)
From Eq.(1) and Eq.(2), it can be concluded that traditional adversarial attacks rely on specify benign images $x$ as the basis for generating adversarial examples $x^{\prime}$.
@@ III-B Diffusion model
Denoising Diffusion Models (DDMs) [33] are a type of generative model that produces images by progressively suppressing noise, starting from pure noise and iterating until a clean image is generated. The process consists of two stages: Forward diffusion and reverse diffusion.
Forward Diffusion Process: In this stage, clean images are corrupted by adding Gaussian noise in a series of steps, eventually transforming the image into pure noise. At each step $t$, the noisy image $x_{t}$ is computed as
$x_{t+1}=\sqrt{a_{t}x_{t}+\sqrt{1-\alpha_{t}\epsilon}},$ (3)
where $a_{t}$ is the noise-scaling factor, and $\epsilon$ represents the Gaussian noise.
Reverse Diffusion Process: Starting from pure noise, the model iteratively denoises the image, generating a series of intermediate images that gradually resemble the original image. The reverse process is modeled as
$p_{\theta}\left(x_{t-1}\mid x_{t}\right)$ $=\mathcal{N}\left(x_{t-1};\mu_{\theta}\left(x_{t},t\right),\Sigma_{\theta}\left(x_{t},t\right)\right),$ (4)
where $\mu_{\theta}$ and $\Sigma_{\theta}$ are the learned parameters for mean and variance, respectively, guiding the denoising process.
To accelerate sampling, Denoising Diffusion Implicit Models (DDIM) [34] introduce a non-stochastic, efficient sampling approach. DDIM avoids fully stochastic sampling, offering a faster alternative by approximating the reverse steps as
$x_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\cdot\mu_{\theta}(x_{t},t)+\sqrt{1-\bar{\alpha}_{t-1}}\cdot\epsilon,$ (5)
which reduces the number of steps required to generate an example, making the model more practical for real-time applications.
@@ III-C EDICT
In recent years, EDICT [35] has been proposed to address the problem of reconstruction and content loss in model-generated images due to error propagation during DDM backpropagation [36]. EDICT mathematically and precisely inverts real- and model-generated images by coupling latent noise variables. Specifically, EDICT utilizes two alternating sequences, $x_{t}$ and $y_{t}$, in the symmetric coupling layer of the DDM. These sequences are used to track and reconstruct the states of latent variables during the generation process. The forward updating of $x_{t}$ and $y_{t}$ during the DDM inverse propagation is then realized through the following coupling steps:
$x_{t}^{inter}$ $=a_{t}\cdot x_{t}+b_{t}\cdot\Theta_{(t,C)}(y_{t});$ (6)
$y_{t}^{inter}$ $=a_{t}\cdot y_{t}+b_{t}\cdot\Theta_{(t,C)}(x_{t}^{inter});$
$x_{t-1}$ $=p\cdot x_{t}^{inter}+(1-p)\cdot y_{t}^{inter};$
$y_{t-1}$ $=p\cdot y_{t}^{inter}+(1-p)\cdot x_{t-1},$
where $a_{t}$ and $b_{t}$ are time-related coefficients, EDICT includes the internal transformations and backward updates of the variables. It generates a new $x_{t-1}$ and $y_{t-1}$ by mixing the internal representations of $x_{t}$ and $y_{t}$ through the averaging parameter $p\in(0,1)$, which is derived using linear equations through the above equation to ensure that the entire diffusion process can be performed exactly in the reverse order:
@@
$x_{t+1}$ $=(x_{t+1}^{inter}-b_{t+1}\cdot\Theta_{(t+1,C)}(y_{t+1}))/a_{t+1}.$